Prove the trigonometric identity: sin(x + y) sin(x - y) = sin’ x – sin? y. Which identity is used to prove it true?
Answer #1

Identities Used are ::

Sin(x+y) = sin x cos y + cos x sin y

sin(x-y) = sin x cos y - cos x sin y

Given, = Sin (x+y) sin(x-Y) (Sinx Cosy + Cosx Siny) (Sinx Cosy - Cosx Siny) Sint cosy. - cosx sinly Sinx (1-siny) - (1- Sinx